OpenStudy (anonymous):

if f(x) = x^2 + 1 and g(x) = 3x + 1 find 2*f(4)

OpenStudy (anonymous):

18 34 65

OpenStudy (mayankdevnani):

\[\large \bf f(x)=x^2+1\] so to find f(4),just plug the value of x=4 you get \[\large \bf f(4)=4^2+1=16+1=17\] but our question says that :- \[\large \bf 2 \times f(4)=2 \times 17=34=\color{red}{Answer}\]

OpenStudy (mayankdevnani):

understood? @smiley2015

OpenStudy (anonymous):

yes

OpenStudy (mayankdevnani):

good :)

OpenStudy (mayankdevnani):

@smiley2015 just click the BEST RESPONSE right side to my comment if you want to give medal

OpenStudy (mayankdevnani):

you did it .... good :)